OpenStudy (anonymous):

Judy works at a local bowling alley. She currently makes $300 per week after taxes plus $60 extra cash every other week for cleaning the lanes. Judy is trying to save up for a car. She really wants a Bug. She knows that she can buy a used Bug in good condition for between $6,600 and $9,900. The equation below represents this situation where x is the number of weeks worked and P is the price of the car: What is the minimum number of weeks that Judy needs to work to buy her Bug? 30 weeks 20 weeks 25 weeks 10 weeks

OpenStudy (anonymous):

$300 per week and 60 every other week is something like $330 every week. if she works \(x\) weeks then she makes \(330x\) dollars.

OpenStudy (anonymous):

the equation below is missing, but it should actually be an inequality. \[6600\leq 330x\leq 9900\]

OpenStudy (anonymous):

i think the answer would be 30?

OpenStudy (anonymous):

to find \(x\) divide across by \(330\) and get \[20\leq x\leq 30\] so she must work AT LEAST 20 hours to get the car